OpenStudy (anonymous):

Evaluate f(x+h)-f(x)/h and simplify if f(x)=x^2-2x

OpenStudy (anonymous):

Ok so first write f(x+h)

OpenStudy (anonymous):

K got it

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

I wrote that down?

OpenStudy (anonymous):

I just wrote down f(x+h) like you said ha

OpenStudy (anonymous):

Oh, no I mean evaluate it. Sorry f(x+h) = ?

OpenStudy (anonymous):

So where there is an x in f(x) put an x+h.

OpenStudy (anonymous):

Okay so it's f(x+h)=(x+h)^2-2(x+h)?

OpenStudy (anonymous):

Right! Now multiply all that out

OpenStudy (anonymous):

So it's -2(x+h)^3?

OpenStudy (anonymous):

Nope..\[f(x+h) = (x+h)^2 - 2(x+h)\] \(=(x+h)(x+h) - 2(x+h)\) \(=(x+h)(x+h) - 2(x+h)\) \(=x^2 + 2xh + h^2 - 2(x+h)\) \(=x^2 + 2xh + h^2 - 2x-2h\)

OpenStudy (anonymous):

You have to foil out the (x+h)^2

OpenStudy (anonymous):

Oh goodness

OpenStudy (anonymous):

It's not as scary as it looks

OpenStudy (anonymous):

So now we want to know what \(f(x+h) - f(x)\) is because we want to divide it by h

OpenStudy (anonymous):

So what is -f(x) ?

OpenStudy (anonymous):

Okay hold on uno momento, I dont' understand where you got the last part of the foiled equation?? How did it turn into x^2+2xh+h^2?

OpenStudy (anonymous):

Ok. \[(x+h)^2\]\[=(x+h)(x+h)\]If you know FOIL you take ther First, Outer, Inner, and Last terms and multiply them, then sum the results. \(=(x\cdot x) + (x\cdot h) + (h \cdot x) + (h\cdot h)\) first outer inner last \[=x^2 + hx + hx + h^2\]\[= x^2 + 2hx + h^2\]

OpenStudy (anonymous):

Okay so it was two different steps? It was (x+h)(x+h)-2(x+h)+(x+h)(x+h)-2(x+h) and then x\[^{2}\]+2xh+h\[^{2}\]=x\[^{2}\]+2xh+h\[^{2}\]-2x-2h?

OpenStudy (anonymous):

Whoa that was weird...

OpenStudy (anonymous):

I'm not sure I follow all that.

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

So it was two different steps? (x+h)(x+h)-2(x+h)=(x+h)(x+h)-2(x+h) and then x^2+2xh+h^2-2(x+h)=x^2+2xh+h^2-2x-2h?

OpenStudy (anonymous):

Yes, those are separate steps

OpenStudy (anonymous):

So now we have an expression for f(x+h) now from that we must subtract f(x)

OpenStudy (anonymous):

So subtract all of the f(x+h) from the equation?

OpenStudy (anonymous):

\[f(x+h) = x^2 + 2xh + h^2 - 2x - 2h\]\[f(x) = x^2 - 2x\] \[\implies f(x+h) - f(x) = [x^2 + 2xh + h^2 - 2x - 2h] - [x^2 - 2x]\] ^This is f(x+h) ^This is f(x)

OpenStudy (anonymous):

So what do we have left after we subtract?

OpenStudy (anonymous):

(2xh+h^2-2h)?

OpenStudy (anonymous):

f(x+h)-F(x)=(2xh+h^2-2h)

OpenStudy (anonymous):

That's exactly right. Sorry someone called.

OpenStudy (anonymous):

So now we divide that by h (cancelling one factor of h from each term)

OpenStudy (anonymous):

Oh it's alright!

OpenStudy (anonymous):

So it's f(x+h)-f(x)=(2x+h-2)?

myininaya (myininaya):

(f(x+h)-f(x))/h=2x+h-2

OpenStudy (anonymous):

So I'm right then myininaya?

myininaya (myininaya):

with that one little correct yes

myininaya (myininaya):

correction*

OpenStudy (anonymous):

Okay but why is it divided by h?

myininaya (myininaya):

you want to find \[\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

Because you had it that way in the original problem

OpenStudy (anonymous):

Right, duh. Sorry my mind is kinda trashed

OpenStudy (anonymous):

We found that \[f(x+h) - f(x) = 2xh + h^2 - 2h\] Therefore \[\frac{f(x+h) - f(x)}{h} = \frac{2xh + h^2 - 2h}{h}=2x + h - 2\]

OpenStudy (anonymous):

Excellent! Thank you so much